y^2 -2x^2 -16=0 Write the equation in standard form. State whether the graph of the equation is a parabola, circle, ellipse, or hyperbola .

Answer by lwsshak3(11628)

y^2 -2x^2 -16=0 Write the equation in standard form. State whether the graph of the equation is a parabola, circle, ellipse, or hyperbola .
y^2-2x^2=16
y^2/16-x^2/8=1
This is an equation of a hyperbola with vertical transverse axis of the standard form:
(y-k)^2/a^2-(x-k)^2/b^2=1, (h,k) being the (x,y) coordinates of the center.